Consider the line y=2/5x-3
Find the equation of the line that is parallel to this line and passes through the point(5, 4).
Find the equation of the line that is perpendicular to this line and passes through the point(5, 4)

Answer by TimothyLamb(4379)

L1: y = 2/5x - 3
---
line parallel to L1 and through point (5, 4)
parallel lines have equal slope:
m = 2/5
---
use the point-slope form:
point = (5, 4)
slope = 2/5 = 0.4
---
to solve the linear equation, plug this:
5 4 0.4
into the "Point-Slope form: x1 y1 m" input box of this solver:
https://sooeet.com/math/linear-equation-solver.php
---
slope = 2/5
y-intercept = 2
---
Answer1:
the slope-intercept form:
y = 2/5x + 2
---
---
---
line perpendicular to L1 and through point (5, 4)
perpendicular lines have negative reciprocal slopes:
m = -5/2 = -2.5
---
use the point-slope form:
point = (5, 4)
slope = -5/2 = -2.5
---
to solve the linear equation, plug this:
5 4 -2.5
into the "Point-Slope form: x1 y1 m" input box of this solver:
https://sooeet.com/math/linear-equation-solver.php
---
slope = -5/2
y-intercept = 16.5
---
Answer2:
the slope-intercept form:
y = -5/2x + 16.5
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php